OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4 What is the resulting equation?

OpenStudy (jackellyn):

2x + (3x +4) = 6 2x + 3x + 4 = 6 5x + 4 = 6

OpenStudy (anonymous):

thx i was just checking to be sure

OpenStudy (jackellyn):

no problem